OpenStudy (anonymous):

For 2x – 3y = –7, when x = 4, What is the value of y?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

first there again

OpenStudy (anonymous):

2x - 3y = -7 When x = 4: 2(4) - 3y = -7 = 8 - 3y = -7 Subtract 8 from each side: -3y = -7 - 8 = -3y = -15 Divide each side by 3: -y = -5 Multiply each side by -1: y = 5 (You could have divided each side by -3 to get y = 5 as well.)